### v2

# On the attainment of the Wasserstein–Cramer–Rao lower bound

Abstract
Recently, a Wasserstein analogue of the Cramer–Rao inequality has been developed using the Wasserstein information matrix (Otto metric). This inequality provides a lower bound on the Wasserstein variance of an estimator, which quantifies its robustness against additive noise. In this study, we investigate conditions for an estimator to attain the Wasserstein–Cramer–Rao lower bound (asymptotically), which we call the (asymptotic) Wasserstein efficiency. We show a condition under which Wasserstein efficient estimators exist for one-parameter statistical models. This condition corresponds to a recently proposed Wasserstein analogue of one-parameter exponential families (e-geodesics). We also show that the Wasserstein estimator, a Wasserstein analogue of the maximum likelihood estimator based on the Wasserstein score function, is asymptotically Wasserstein efficient in location-scale families.
1 Introduction
The Cramer–Rao inequality is a well-known classical theorem in statistics. It provides a lower bound on the variance of (unbiased) estimators through the inverse of the Fisher information matrix. An estimator is said to be (asymptotically) Fisher efficient if it attains the Cramer–Rao lower bound (asymptotically). For one-parameter statistical models, an estimator is Fisher efficient if and only if the model is an exponential family and it is the maximum likelihood estimator (MLE) of its expectation parameter [10, 16]. For general models, the MLE is asymptotically Fisher efficient under regularity conditions [14]. In information geometry, the Fisher information is adopted as a Riemannian metric on the parameter space and is closely connected to the Kullback–Leibler divergence [2].
The Wasserstein distance is defined as the optimal transport cost between probability distributions and it induces another geometric structure on the space of probability distributions [15]. The Wasserstein geometry has been widely applied in many fields, including statistics and machine learning [5, 13]. Recently, Li and Zhao [11] developed Wasserstein counterparts of information geometric concepts such as the Wasserstein information matrix and Wasserstein score function. They also derived the Wasserstein–Cramer–Rao inequality, which gives a lower bound on the Wasserstein variance of an estimator by the inverse of the Wasserstein information matrix. Whereas the usual variance quantifies the accuracy of an estimator, the Wasserstein variance can be interpreted as the robustness of an estimator against additive noise [1]. Li and Zhao [11] also proposed an estimator called the Wasserstein estimator as the zero point of the Wasserstein score function.
In this study, we investigate conditions for an estimator to attain the Wasserstein–Cramer–Rao lower bound (asymptotically), which we call the (asymptotic) Wasserstein efficiency. In Section 2, we briefly review the Wasserstein–Cramer–Rao inequality. In Section 3, we focus on one-parameter models and derive a condition for Wasserstein efficiency in finite samples, which corresponds to recently proposed Wasserstein analogue of one-parameter exponential families (e-geodesics) [3]. In Section 4, we focus on location-scale families and show that the Wasserstein estimator is asymptotically Wasserstein efficient.
2 Wasserstein–Cramer–Rao Inequality
In this section, we briefly review the Wasserstein information matrix and Wasserstein–Cramer–Rao inequality introduced by Li and Zhao [11]. We consider a parametric density $p(x;\theta)$ on $\mathbb{R}^{d}$ with parameter $\theta\in\mathbb{R}^{p}$ in the following.
The Wasserstein score function $\Phi_{i}(x;\theta)$ for $i=1,\dots,p$ is defined as the solution to the partial differential equation
$\frac{\partial}{\partial\theta_{i}}p(x;\theta)+\nabla_{x}\cdot\left(p(x;\theta)\nabla_{x}\Phi_{i}(x;\theta)\right)=0$ (1)
satisfying ${\rm E}_{\theta}[\Phi_{i}(x;\theta)]=0$, where $\nabla_{x}\cdot f$ is the divergence of a vector field $f=(f_{1},\dots,f_{d})$ given by
$\nabla_{x}\cdot f=\sum_{i=1}^{d}\frac{\partial f_{i}}{\partial x_{i}}$
and $\nabla_{x}g$ is the gradient of a function $g$ given by
$\nabla_{x}g=\left(\frac{\partial g}{\partial x_{1}},\dots,\frac{\partial g}{\partial x_{d}}\right)^{\top}.$
Note that (1) is often called the continuity equation in the dynamic formulation of the optimal transport problem [4, 15]. In this context, the Wasserstein score function $\Phi_{i}(x;\theta)$ can be viewed as a solution of the Hamilton–Jacobi equation (up to additive constant), where $\theta_{i}$ is adopted as the time variable.
The Wasserstein information matrix $G_{W}(\theta)\in\mathbb{R}^{d\times d}$ is defined as
$G_{W}(\theta)_{ij}$ $={\rm E}_{\theta}\left[(\nabla_{x}\Phi_{i}(x;\theta))^{\top}(\nabla_{x}\Phi_{j}(x;\theta))\right],\quad i,j=1,\dots,d.$
Recall that the $L^{2}$-Wasserstein distance $W_{2}(p,q)$ between two probability densities $p$ and $q$ on $\mathbb{R}^{d}$ is defined by
$W_{2}(p,q)=\frac{1}{2}\inf_{X,Y}\ {\rm E}[\|X-Y\|^{2}]^{\frac{1}{2}},$
where the infimum is taken over all joint distributions (coupling) of $(X,Y)$ with marginal distributions of $X$ and $Y$ equal to $p$ and $q$, respectively. The Wasserstein information matrix appears in the quadratic approximation of the $L^{2}$ Wasserstein distance:
$W_{2}(p_{\theta},p_{\theta+\Delta\theta})^{2}=\frac{1}{2}\Delta\theta^{\top}G_{W}(\theta)\Delta\theta+o(\|\Delta\theta\|^{2}).$
For a statistic $a(x)\in\mathbb{R}^{l}$, its Wasserstein variance ${\rm Var}_{\theta}^{W}[a(x)]\in\mathbb{R}^{l\times l}$ is defined by
${\rm Var}_{\theta}^{W}[a(x)]_{ij}={\rm E}_{\theta}\left[(\nabla_{x}a_{i}(x))^{\top}(\nabla_{x}a_{j}(x))\right],\quad i,j=1,\dots,l.$
Note that the Wasserstein information matrix is the Wasserstein variance of the Wasserstein score function.
Lemma 1 (Wasserstein–Cramer–Rao inequality [11]).
For a statistic $a(x)\in\mathbb{R}^{l}$,
${\rm Var}_{\theta}^{W}[a(x)]\succeq\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[a(x)]\right)^{\top}G_{W}(\theta)^{-1}\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[a(x)]\right),$ (2)
where
$\frac{\partial}{\partial\theta}{\rm E}_{\theta}[a(x)]:=\left(\frac{\partial}{\partial\theta_{i}}{\rm E}_{\theta}[a_{j}(x)]\right)_{ij}\in\mathbb{R}^{d\times l}.$
In particular, if $d=l$ and $a(x)$ is an unbiased estimator of $\theta$ (${\rm E}_{\theta}[a(x)]=\theta$), then
${\rm Var}_{\theta}^{W}[a(x)]\succeq G_{W}(\theta)^{-1}.$
We refer to the inequality (2) as the Wasserstein–Cramer–Rao inequality in the following. Recently, [1] discussed its connection to robustness of an estimator against additive noise. We also note that the Wasserstein–Cramer–Rao inequality has been obtained independently in statistical physics and called the short-time thermodynamic uncertainty relation [6, 7, 9, 12].
In this paper, we say that an estimator is (asymptotically) Wasserstein efficient if it attains the Wasserstein–Cramer–Rao lower bound (asymptotically). We investigate conditions of (asymptotic) Wasserstein efficiency in the following.
3 Wasserstein efficiency in one-parameter models
In this section, we focus on scalar estimators for one-parameter models $p(x;\theta)$ on $\mathbb{R}^{d}$ (i.e., $l=p=1$). In this setting, attainment of the (original) Cramer–Rao lower bound has been studied well [16, 10]. Namely, a scalar estimator $a(x)$ attains the Cramer–Rao lower bound for every $\theta$ if and only if the model is a one-parameter exponential family
$p(x;\theta)=g(x)\exp(\theta T(x)-\psi(\theta))$
and the estimator $a(x)$ is the maximum likelihood estimator of its expectation parameter $T(x)$ (or its affine transform). Note that a one-parameter exponential families corresponds to an e-geodesic with respect to the Fisher metric in information geometry [2].
Now, we consider the Wasserstein case. We write $\Phi(x;\theta)=\Phi_{1}(x;\theta)$ for convenience. Since the Wasserstein–Cramer–Rao inequality (2) is derived from the Cauchy–Schwarz inequality [11, 1], its equality condition is obtained as follows.
Theorem 1.
Let $p(x;\theta)$ be a one-parameter model on $\mathbb{R}^{d}$ and $a(x)$ be a scalar estimator. Then,
${\rm Var}_{\theta}^{W}[a(x)]=\frac{1}{G_{W}(\theta)}\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[a(x)]\right)^{2}$
holds if and only if
$a(x)=u(\theta)\Phi(x;\theta)+v(\theta)$
for some $u(\theta)$ and $v(\theta)$.
Proof.
For random vectors $U$ and $V$,
$0\leq{\rm E}\|tU+V\|^{2}={\rm E}[\|U\|^{2}]t^{2}+2{\rm E}[U^{\top}V]t+{\rm E}[\|V\|^{2}]$
for every $t$. Thus, by considering the discriminant of the quadratic equation,
${\rm E}[U^{\top}V]^{2}\leq{\rm E}[\|U\|^{2}]{\rm E}[\|V\|^{2}],$ (3)
where the equality holds if and only if $U$ and $V$ are linearly dependent.
From the definition of the Wasserstein score function,
$\frac{\partial}{\partial\theta_{i}}{\rm E}_{\theta}[a(x)]$ $=\int a(x)\frac{\partial}{\partial\theta_{i}}p(x;\theta){\rm d}x$
$=-\int a(x)\nabla_{x}\cdot(p(x;\theta)\nabla_{x}\Phi(x;\theta)){\rm d}x$
$=-\int(\nabla_{x}\cdot(a(x)p(x;\theta)\nabla_{{x}}\Phi(x;\theta))-(\nabla_{{x}}a(x))^{\top}(\nabla_{{x}}\Phi(x;\theta))p(x;\theta)){\rm d}x$
$={\rm E}_{\theta}[(\nabla_{{x}}a(x))^{\top}(\nabla_{{x}}\Phi(x;\theta))].$
Also,
${\rm E}_{\theta}[\|\nabla_{x}a(x)\|^{2}]={\rm Var}_{\theta}^{W}[a(x)],\quad{\rm E}_{\theta}[\|\nabla_{x}\Phi(x;\theta)\|^{2}]=G_{W}(\theta).$
Thus, the inequality (3) with $U=\nabla_{x}a(x)$ and $V=\nabla_{x}\Phi(x;\theta)$ is equivalent to the Wasserstien–Cramer–Rao inequality (2):
${\rm Var}_{\theta}^{W}[a(x)]\geq\frac{1}{G_{W}(\theta)}\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[a(x)]\right)^{2}.$
Therefore, the Wasserstein–Cramer–Rao lower bound is attained if and only if $\nabla_{x}a(x)$ and $\nabla_{x}\Phi(x;\theta)$ are linearly dependent. This condition is rewritten as $a(x)=u(\theta)\Phi(x;\theta)+v(\theta)$ for some $u(\theta)$ and $v(\theta)$. ∎
Recently, [3] provided a framework of Wasserstein information geometry, where the e-connection is defined as the dual of the m-connection with respect to the Wasserstein metric (Otto metric). This e-connection is different from the one in the usual information geometry based on the Fisher metric [2]. In particular, [3] showed that the e-geodesics are characterized as one-parameter models with fixed Wasserstein score functions: $\Phi(x;\theta)\equiv\Phi(x)$. Note that it is different from the displacement interpolation [15], which corresponds to the geodesic with respect to the Levi-Civita connection. From this viewpoint, Theorem 1 can be rewritten as follows.
Corollary 1.
For a one-parameter model $p(x;\theta)$ on $\mathbb{R}^{d}$, a scalar estimator $a(x)$ attains the Wasserstein–Cramer–Rao lower bound for every $\theta$ if and only if the model corresponds to an e-geodesic with respect to the Wasserstein metric (up to monotone reparametrization) and $a(x)$ is an affine function of its Wasserstein score function.
Proof.
From Theorem 1, an estimator $a(x)$ attains the Wasserstein–Cramer–Rao lower bound for every $\theta$ if and only if $a(x)=u(\theta)\Phi(x;\theta)+v(\theta)$ for every $\theta$. In this case, the Wasserstein score functions $\Phi(x;\theta_{1})$ and $\Phi(x;\theta_{2})$ are affinely related for every $\theta_{1}$ and $\theta_{2}$. Thus, after some monotone reparametrization $\widetilde{\theta}=h(\theta)$, the Wasserstein score function $\Phi(x;\widetilde{\theta})$ becomes the same for every $\widetilde{\theta}$. It is similar to a unit-speed parametrization of a curve on a Riemannian manifold. Therefore, the model $p(x;\theta)$ is an e-geodesic with respect to the Wasserstein metric introduced in [3] up to monotone reparametrization. ∎
Since an e-geodesic with respect to the Wasserstein metric can be viewed as a Wasserstein analogue of the one-parameter exponential family, Corollary 1 is a natural generalization of the classical result on attainment of the Cramer–Rao lower bound [16, 10]. We give several examples of Wasserstein efficient estimators.
Proposition 1.
  1. 1.
    The sample mean
    $a(x)=\frac{1}{n}\sum_{t=1}^{n}x_{t}$
    is Wasserstein efficient if and only if the model is the location family
    $p(x;\theta)=f(x-\theta).$
  2. 2.
    The sample mean of the square
    $a(x)=\frac{1}{n}\sum_{t=1}^{n}x_{t}^{2}$
    is Wasserstein efficient if and only if the model is the scale family
    $p(x;\theta)=\frac{1}{\theta}f\left(\frac{x}{\theta}\right).$
4 Wasserstein efficiency in location-scale families
In this section, we consider location-scale families on $\mathbb{R}$:
$p(x;\theta)=\frac{1}{\sigma}f\left(\frac{x-\mu}{\sigma}\right),\quad\theta=(\mu,\sigma),$ (4)
where $f$ is a probability density on $\mathbb{R}$ with mean zero and variance one (e.g., ${\rm N}(0,1)$). The mean and variance of $p(x;\theta)$ are $\mu$ and $\sigma^{2}$, respectively. Its Wasserstein score function is
$\Phi_{\mu}(x;\theta)=x-\mu,\quad\Phi_{\sigma}(x;\theta)=\frac{(x-\mu)^{2}}{2\sigma}-\frac{\sigma}{2}.$ (5)
Suppose that we have $n$ independent observations $x_{1},\dots,x_{n}$ from $p(x;\theta)$. Then, the Wasserstein estimator $\hat{\theta}_{W}=(\hat{\mu}_{W},\hat{\sigma}_{W})$ is defined as the zero point of the Wasserstein score function [11]:
$\sum_{t=1}^{n}\Phi_{\mu}(x_{t};\hat{\theta}_{W})=\sum_{t=1}^{n}\Phi_{\sigma}(x_{t};\hat{\theta}_{W})=0.$
From (5), it is given by the sample mean and sample standard deviation:
$\hat{\mu}_{W}=\bar{x}=\frac{1}{n}\sum_{t=1}^{n}x_{t},\quad\hat{\sigma}_{W}=\sqrt{\frac{1}{n}\sum_{t=1}^{n}(x_{t}-\bar{x})^{2}}.$ (6)
Theorem 2.
For the location-scale family (4), the Wasserstein estimator $\hat{\theta}_{W}=(\hat{\mu}_{W},\hat{\sigma}_{W})$ in (6) asymptotically attains the Wasserstein–Cramer–Rao lower bound:
$n\left({\rm Var}_{\theta}^{W}[\hat{\theta}_{W}]-\frac{1}{n}\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[\hat{\theta}_{W}]\right)^{\top}G_{W}(\theta)^{-1}\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[\hat{\theta}_{W}]\right)\right)\to O$
as $n\to\infty$.
Proof.
From (6),
$\nabla\hat{\mu}_{W}=\frac{1}{n}\left(1,\dots,1\right),\quad\nabla\hat{\sigma}_{W}=\frac{1}{n\hat{\sigma}_{W}}(x_{1}-\bar{x},\dots,x_{n}-\bar{x}).$
Thus,
${\rm Var}_{\theta}^{W}[\hat{\theta}_{W}]=\begin{pmatrix}{\rm E}_{\theta}[\nabla\hat{\mu}_{W}\cdot\nabla\hat{\mu}_{W}]&{\rm E}_{\theta}[\nabla\hat{\mu}_{W}\cdot\nabla\hat{\sigma}_{W}]\\ {\rm E}_{\theta}[\nabla\hat{\sigma}_{W}\cdot\nabla\hat{\mu}_{W}]&{\rm E}_{\theta}[\nabla\hat{\sigma}_{W}\cdot\nabla\hat{\sigma}_{W}]\end{pmatrix}=\frac{1}{n}\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
On the other hand,
$G_{W}(\theta)=n\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad{\rm E}_{\theta}[\hat{\theta}_{W}]=\begin{pmatrix}\mu\\ c_{n}\sigma\end{pmatrix},$
where $c_{n}$ is the expected value of the sample standard deviation of $x_{1},\dots,x_{n}\sim f$. Thus,
$\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[\hat{\theta}_{W}]\right)^{\top}G_{W}(\theta)^{-1}\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[\hat{\theta}_{W}]\right)=\frac{1}{n}\begin{pmatrix}1&0\\ 0&c_{n}^{2}\end{pmatrix}.$
From the law of large numbers and continuous mapping theorem,
[…]
${\rm E}[\hat{\sigma}_{W}1(\hat{\sigma}_{W}>M)]\leq\frac{{\rm E}[\hat{\sigma}_{W}^{2}]}{M}=\frac{\sigma^{2}}{M},$
shows the uniform integrability of $\hat{\sigma}_{W}$: $\sup_{n}{\rm E}[\hat{\sigma}_{W}1(\hat{\sigma}_{W}>M)]\to 0$ as $M\to\infty$. Therefore, we have ${\rm E}[\hat{\sigma}_{W}]\to\sigma$ and thus $c_{n}\to 1$ as $n\to\infty$. Hence,
$n\left({\rm Var}_{\theta}^{W}[\hat{\theta}_{W}]-\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[\hat{\theta}_{W}]\right)^{\top}G_{W}(\theta)^{-1}\left(\frac{\partial}{\partial\theta}{\rm E}_{\theta}[\hat{\theta}_{W}]\right)\right)=\begin{pmatrix}0&0\\ 0&1-c_{n}^{2}\end{pmatrix}\to O$
as $n\to\infty$. ∎
Since the Wasserstein variance of an estimator quantifies its robustness in terms of the increase of its variance due to noise contamination [1], Theorem 2 implies that the Wasserstein estimator is robust against additive noise in location-scale families. We confirm this for the Laplace distribution:
$p(x;\theta)=\frac{1}{\sqrt{2}\sigma}\exp\left(-\sqrt{2}\left|\frac{x-\mu}{\sigma}\right|\right).$ (7)
For $\mu$, the Wasserstein estimator $\hat{\mu}_{\mathrm{W}}$ is the sample mean while the MLE $\hat{\mu}_{\mathrm{ML}}$ is the sample median. Both estimators are unbiased and their variances are
$\mathrm{Var}_{\theta}(\hat{\mu}_{W})=\frac{\sigma^{2}}{n},\quad\mathrm{Var}_{\theta}(\hat{\mu}_{\mathrm{ML}})=\frac{\sigma^{2}}{2n}+o\left(\frac{1}{n}\right).$
Now, suppose that we have noisy observations $\tilde{x}_{1}=x_{1}+z_{1},\dots,\tilde{x}_{n}=x_{n}+z_{n}$ instead of $x_{1},\dots,x_{n}$, where $z_{1},\dots,z_{n}\sim{\rm N}(0,\varepsilon^{2})$ are independent Gaussian noise with variance $\varepsilon^{2}$. Then,
$\frac{\mathrm{Var}_{\theta}(\hat{\mu}_{\mathrm{W}}(x_{1}+z_{1},\dots,x_{n}+z_{n}))-\mathrm{Var}_{\theta}(\hat{\mu}_{\mathrm{W}}(x_{1},\dots,x_{n}))}{\varepsilon^{2}}=\frac{1}{n},$
which does not depend on $\varepsilon^{2}$. On the other hand, as shown in Appendix,
$\frac{\mathrm{Var}_{\theta}(\hat{\mu}_{\mathrm{ML}}(x_{1}+z_{1},\dots,x_{n}+z_{n}))-\mathrm{Var}_{\theta}(\hat{\mu}_{\mathrm{ML}}(x_{1},\dots,x_{n}))}{\varepsilon^{2}}\approx\frac{2\sigma}{\sqrt{\pi}n\varepsilon}$ (8)
for large $n$, which diverges as $\varepsilon^{2}\to 0$. Therefore, the Wasserstein estimator is more robust than MLE against small noise.
Appendix A Derivation of (8)
From Section 13 of [8], the asymptotic distribution of the sample median of $n$ independent samples $x_{1},\dots,x_{n}\sim p$ is
$\sqrt{n}({\rm median}(x_{1},\dots,x_{n})-m)\to{\rm N}\left(0,\frac{1}{4p(m)^{2}}\right),$
as $n\to\infty$, where $m$ is the median of $p$. Thus,
${\rm Var}({\rm median}(x_{1},\dots,x_{n}))\approx\frac{1}{4np(m)^{2}}$ (9)
for large $n$. Therefore, for the Laplace distribution (7),
$\mathrm{Var}_{\theta}(\hat{\mu}_{\mathrm{ML}}(x_{1},\dots,x_{n}))\approx\frac{\sigma^{2}}{2n}$ (10)
for large $n$. On the other hand, the probability density of the noisy observation $\tilde{x}=x+z$ with $z\sim{\rm N}(0,\varepsilon^{2})$ is given by the convolution
$p(\tilde{x})$ $=\int_{-\infty}^{\infty}\frac{1}{\sqrt{2}\sigma}\exp\left(-\sqrt{2}\left|\frac{\tilde{x}-z-\mu}{\sigma}\right|\right)\frac{1}{\sqrt{2\pi\varepsilon^{2}}}\exp\left(-\frac{z^{2}}{2\varepsilon^{2}}\right){\rm d}z.$
At the median $\tilde{x}=\mu$,
$p(\tilde{x}=\mu)$ $=\int_{-\infty}^{\infty}\frac{1}{\sqrt{2}\sigma}\exp\left(-\sqrt{2}\left|\frac{-z}{\sigma}\right|\right)\frac{1}{\sqrt{2\pi\varepsilon^{2}}}\exp\left(-\frac{z^{2}}{2\varepsilon^{2}}\right){\rm d}z$
$=2\int_{0}^{\infty}\frac{1}{2\sqrt{\pi}\sigma\varepsilon}\exp\left(-\sqrt{2}\frac{z}{\sigma}-\frac{z^{2}}{2\varepsilon^{2}}\right){\rm d}z$
$=\frac{1}{\sqrt{\pi}\sigma\varepsilon}\int_{0}^{\infty}\exp\left(-\frac{1}{2\varepsilon^{2}}\left(z+\frac{\sqrt{2}\varepsilon^{2}}{\sigma}\right)^{2}+\frac{\varepsilon^{2}}{\sigma^{2}}\right){\rm d}z$
$=\frac{\sqrt{2}}{\sigma}\Phi\left(-\frac{\sqrt{2}\varepsilon}{\sigma}\right)\exp\left(\frac{\varepsilon^{2}}{\sigma^{2}}\right)$
$=\frac{1}{\sqrt{2}\sigma}-\frac{\sqrt{2}}{\sqrt{\pi}\sigma^{2}}\varepsilon+O(\varepsilon^{2})$
as $\varepsilon\to 0$, where $\Phi$ is the cumulative distribution function of the standard Gaussian ${\rm N}(0,1)$. Thus, by using (9),
${\rm Var}_{\theta}(\hat{\mu}_{\mathrm{ML}}(x_{1}+z_{1},\dots,x_{n}+z_{n}))\approx\frac{\sigma^{2}}{2n}\left(1+\frac{4}{\sqrt{\pi}\sigma}\varepsilon+O(\varepsilon^{2})\right)$ (11)
for large $n$. Combining (10) and (11) yields (8).
Acknowledgments
We thank Keiya Sakabe and Sosuke Ito for helpful comments. Takeru Matsuda was supported by JSPS KAKENHI Grant Numbers 19K20220, 21H05205, 22K17865 and JST Moonshot Grant Number JPMJMS2024.